Suppose you know that a companys slock currenty sets for $66 40 per share and the r its the companys policy to always maintain a constant growth rate in its dividends, what is the curnent dividen owidend per shareS atso ?ow that the los resam on the stock is eventy dvided between captat gans ned and dvidend yeid calculations and round your answer to 2 decimal places, .g.32.14) Leaming Objective: or 01 Assess how slock prices depend on uture dividends and dvidend growth
Answer #1

Solution:-

Dividend per share $ 3.17 1%

Explanation:-

We know the stock has a required return of 10 percent, and the dividend and capital gains yield are equal, so:

Dividend yield = 1/2(.10)

Dividend yield = .050 = Capital gains yield

Now we know both the dividend yield and capital gains yield. The dividend is simply the stock price times the dividend yield, so:

D1= .050($66.40)

D1= $3.32

This is the dividend next year. The question asks for the dividend this year. Using the relationship between the dividend this year and the dividend next year:

D1= D0(1 + g)

We can solve for the dividend that was just paid:

$3.32 = D0(1 + .050)

D0= $3.32 / 1.050

D0= $3.17

Please Rate or comment if you have any doubt regarding this solution.